OpenStudy (petlovers123):

Netan surveyed 460 teenagers to find out their favorite type of movie. He found that 45% of the teenagers surveyed like comedies. Part A: How many teenagers surveyed like comedies? Show your work. (4 points) Part B: In another survey, Hannah found that 50% of the total number of teenagers surveyed like thrillers. If 40 teenagers like thrillers, what is the total number of teenagers that Hannah surveyed? Show your work. (4 points) Part C: Give an example of a situation in which a percentage is greater than 100%. (2 points) @pooja195 I dont get this at all

OpenStudy (caozeyuan):

total number of objects * percentage of category A= number of objects belong to category A

OpenStudy (petlovers123):

Don't get it

OpenStudy (caozeyuan):

Ok, lets break this down. SO you have a total of 450 students, so your total number is 450, and it says 40% like comedy, so your percentage is 45%, now to get the number of students who like comdey, you just work out 450*45%

OpenStudy (petlovers123):

to get part B

OpenStudy (petlovers123):

or for the whole question

OpenStudy (caozeyuan):

Well, you just do division , beucase now the unknown is the total amount, not the amount of a particular category

OpenStudy (petlovers123):

divide what?

OpenStudy (caozeyuan):

Well, if A*B=C, now you know B and C, how can you get A?

OpenStudy (petlovers123):

A is i think 260

OpenStudy (caozeyuan):

Should be80, right?

OpenStudy (caozeyuan):

cause you have A*50%=40, A=80

OpenStudy (petlovers123):

I'm just very confused this is so hard feel like giving up

OpenStudy (caozeyuan):

Yep, percentage is hard to deal with for a newcomer, reduce it to fractions if you hate it

OpenStudy (petlovers123):

how

OpenStudy (caozeyuan):

for example, 50% is 1/2 because 50%=50/100

OpenStudy (petlovers123):

right

OpenStudy (caozeyuan):

so part B is asking what is the number that will give you 40 if you half it

OpenStudy (caozeyuan):

feel better now?

OpenStudy (petlovers123):

yea

OpenStudy (caozeyuan):

so the answer is obvious, isn't it?

OpenStudy (caozeyuan):

the magic equation to work with is A*B=C, where A is the total number of things, this is the number of students in your question. B is the percentage of a certain category, this is 45% for your part A. C is the number of things within that category. This is the answer for your part A

OpenStudy (caozeyuan):

for part B, because A is the unknown, your do C/B, that gives your 80 as the asnwer

OpenStudy (petlovers123):

sob the answer is 80

OpenStudy (caozeyuan):

yes,and for C you can make up whatever you what, e.g, the weight gain of my pet cat is 110%

OpenStudy (petlovers123):

but for A I'm not aloud to put that for an unknown reason

OpenStudy (caozeyuan):

I don't understnad, is the answer wrong or someting?

OpenStudy (petlovers123):

i don't know when i typed it it said invalid entry

OpenStudy (petlovers123):

I'll figure A out but thnx for all your help